Find the optimum level of output and profit from the cost function

TC = 50 + 6Q2

 and price

P = 100 – 4Q

Also derive marginal cost and marginal revenue. 


1
Expert's answer
2021-04-29T10:23:10-0400

"P=100-4Q
, total revenue =price \times quantity"

"total\space revenue(TR)=(100-4Q)Q=100Q-4Q^2"

"dTR\/dQ=100-8Q=MR"

"MC=dTC\/dQ=12Q"

Profit is maximized at the level of output where MR=MC

"MR=MC"

"100-8Q=12Q"

"20Q=100"

"Q=5 \space (optimal \space output)"

"TR=100Q-4Q^2=100(5)-4(5^2)=400"

"TC=50+6Q^2=50+6(5)^2=200"

"profit=TR-TC=400-200=200 \space (profit\space at\space the \space optimal\space level\space of\space output)"